A 0.86 L sample of nitrogen gas is at a temperature of 45.0°C. If the volume decreases to 0.56 L what is the temperature of the gas?
1
Expert's answer
2019-09-19T05:11:26-0400

"V_1\/T_1=V_2\/T_2"

"T_2=\frac{V_2·T_1}{V_1}=\frac{0.56·(273.15+45.0)}{0.86}=" 207.2K=-66.0C

Answer:-66.0C